# Adversarial Sample-Based Approach for Tighter Privacy Auditing in Final Model-Only Scenarios

### v1

@@ 6 Conclusion
We introduce a novel method for crafting worst-case adversarial samples to achieve tighter lower bounds in privacy auditing of differentially private models, focusing on the final model setting. Unlike traditional canary-based approaches, our technique maximizes the distinguishability between output distributions by designing specialized loss functions, particularly the adaptive loss function $\mathcal{L}_{ade}$. This approach enhances the separability of overlapping regions in the loss distributions, leading to more effective auditing even when only the final model is accessible.
Our experiments on MNIST and CIFAR-10 datasets demonstrate that our method consistently provides tighter empirical lower bounds on privacy leakage compared to baseline methods, both in white-box and black-box settings. For instance, on CIFAR-10 with a theoretical privacy budget of $\varepsilon=10.0$, the traditional canary-based method yields an empirical lower bound of only $0.30$, whereas our approach achieves a significantly tighter lower bound of $4.10$. Notably, our approach is effective even when using in-distribution (ID) samples as canaries, eliminating the need for out-of-distribution (OOD) samples. This robustness across various canary types highlights the practical applicability of our method.
By enabling more precise privacy auditing without relying on specific canary samples or additional assumptions, our work significantly advances the evaluation of differential privacy guarantees in the final model setting.

### v2

@@ 6 Conclusion
We propose a novel method for crafting worst-case adversarial samples that tighten lower bounds in privacy auditing of differentially private models, focusing on the final model setting without additional assumptions. In contrast to traditional canary-based approaches, our method generates adversarial samples that maximize the distinguishability between output distributions. Experiments on the MNIST dataset demonstrate that this approach provides substantially tighter empirical lower bounds on privacy leakage compared to baseline methods. These findings highlight the promise of employing specialized adversarial samples for more rigorous privacy auditing, and pave the way for extending such techniques to more complex datasets and model architectures.